OpenStudy (anonymous):

An equilateral triangle is inscribed in a circle of radius 10 cm. Find: (a) the area of the triangle. (b) the area shaded.

OpenStudy (anonymous):

Can someone help?

OpenStudy (anonymous):

its easy

OpenStudy (anonymous):

equilateral triangle's all sides are equal and angles are equal

OpenStudy (anonymous):

 simple geometry

OpenStudy (anonymous):

use "r" to find the height of the triangle

OpenStudy (anonymous):

or there is a formula you can find a by radius that's r =(root 3/3) *a

OpenStudy (anonymous):

\[r = (\sqrt{3}/3)/a\]

OpenStudy (anonymous):

no. r is not the height. r = (2/3)*h

OpenStudy (anonymous):

\[A =(\sqrt{3}/4) x a ^{2}\] that's for area

OpenStudy (anonymous):

sorry that's root 3 by 4 multiplied by a square

OpenStudy (anonymous):

Haha thanks guys